# Virchow 2: Scaling Self-Supervised Mixed Magnification Models in Pathology

### v1

@@ Abstract
Foundation models are rapidly being developed for computational pathology applications. However, it remains an open question which factors are most important for downstream performance with data scale and diversity, model size, and training algorithm all playing a role. In this work, we present the result of scaling both data and model size, surpassing previous studies in both dimensions, and introduce two new models: Virchow 2, a 632M parameter vision transformer, and Virchow 2G, a 1.85B parameter vision transformer, each trained with 3.1M histopathology whole slide images. To support this scale, we propose domain-inspired adaptations to the DINOv2 training algorithm, which is quickly becoming the default method in self-supervised learning for computational pathology. We achieve state of the art performance on twelve tile-level tasks, as compared to the top performing competing models. Our results suggest that data diversity and domain-specific training can outperform models that only scale in the number of parameters, but, on average, performance benefits from domain-tailoring, data scale, and model scale.
@@ 1 Introduction
Figure 1: The average weighted F1 score across 8 tile-level benchmarks using public data is shown with respect to two scaling axes: model size (left), in terms of the number of parameters, and dataset size (right), in terms of the number of whole slide images (WSIs). Virchow 2 improves over Virchow with pathology-specific modifications to the training method while scaling up the dataset from 1.5M to 3.1M WSIs, with increased diversity. Virchow 2G then scales the model size from 632M to 1.8B parameters (ViT-H to ViT-G). Benchmark performance across foundation models in pathology appears to scale with model and dataset size.
In this work, we explore scaling along both the data and model size axes, while adjusting the training method with domain-inspired modifications. We present Virchow 2 and Virchow 2G, mixed magnification foundation models for computational pathology. Virchow 2, a 632M parameter vision transformer (ViT-H) expands on our work developing Virchow [50], scaling the dataset size from 1.5M WSIs from a single institution with routine hemotoxylin and eosin (H&E) staining, to 3.1M WSIs from globally diverse institutions with diverse staining such as immunohistochemistry (IHC). Virchow 2G builds on Virchow 2 by increasing the model size to 1.8B parameters (ViT-G).
@@ 2.2 CPath Foundation Models
3 Learning in the pathology domain
@@ 3.1 Domain-specific augmentations
Figure 2: a. Virchow 2 and 2G are trained with mixed magnification samples; slides are tiled at 40x, 20x, 10x, or 5x. b. Illustration of the difference between standard crop-and-resize augmentation (top) and the proposed extended-context translation (ECT) augmentation (bottom). ECT is expected to be more suitable for pathology images as it avoids distorting the cell morphology. We tuned the ECT approach to retain the average overlap between crops (shaded, teal). c. A diversity regularization objective seeks to push apart tile embeddings on a hypersphere (top). For Virchow 2 and 2G, we replace the unstable KoLeo regularizer used by Virchow with a kernel density estimator (KDE) that does not produce exploding gradients for similar tiles (bottom), which is particularly important for pathology whole slide images as they contain many highly correlated tiles.
@@ 4 Stabilizing self-supervised vision transformers at scale
Optimizing ViTs requires adaptive methods like AdamW and its variations [Balestriero2023ACO]. The stability of AdamW depends on the efficacy of the warm-up phase, on the learning rate, and on the selection of momentum parameters. Gradient spikes have been correlated with the ratio of first and second moments for an update step which provide insights on when a collapse event is likely to occur [38, 56]. For example, instability has been observed in the weights of the patch embedding layer [16, 56]. It is possible to reduce these issues by lowering the second moment hyperparameter but this may incur reduced downstream task performance. StableAdamW is explored as a means of stabilizing weight updates when the ratio between the first and second moments are too large (such as when nearly converged weights suddenly need a correction during a long training run) by mitigating large update steps with root-mean-square clipping [56].
@@ 5.3 Tile benchmarks
PanMSK XX×$\times$ is an in-distribution cancer detection task performed at 5$\times$, 10$\times$, and 20$\times$ magnifications. 1,196,171 $224\times 224$ pixel tile samples, sourced from 3,999 WSI from MSKCC representing 17 tissue types, are labeled either cancer or benign [50].
@@ 5.5 Ablation results
Table 1: Weighted F1 results for seven evaluation tasks and five model settings with a ViT-B/16. For each change, a checkmark indicates a deviation from the standard DINOv2 training protocol. In-distribution evaluation tasks include PanMSK XX, while out-of-distribution tasks include PCam, CRC, MHIST, and MIDOG. Boldface indicates the best performance and italics indicates the second best performance. In a majority of tasks and on average, ECT, KDE, and no solarization regularization performs best.
6 Results
@@ 6.1 Training data
Figure 3: Overivew of the training dataset. a. A breakdown of the scale of the training dataset with 3.1M whole slide images (WSIs), showing the distribution of slides with hemotoxylin and eosin (H&E) and immunohistochemical (IHC) stains. b. The data are globally diverse, with 15% of the slides and 57% of the patients sourced from many institutions across all continents. These slides were submitted to MSKCC for review. c. The distribution of most severe diagnoses of the WSIs, in decreasing order of severity: cancer, precursor, benign, unknown. d. The top identified tissue types shown as a fraction of the WSIs. There are over 150 additional tissue types in the long tail, summarized here as ‘Other Tissues’ (shown gray, along with ‘Unknown Tissue’, on the right). Online data balancing when training Virchow 2G affected these tissue distributions as shown by the magenta dotted line.
@@ 6.2 Pretraining methodology
Following the modifications made to the ViT-g architecture to maximize compute efficiency [42], we introduce Virchow 2G, a ViT-G/14 that adopts similar architectural changes. Virchow 2G increases the embedding dimension from 1,664 to 1,792 to be a multiple of 256 and increases the number of attention heads from 16 to 28 in order to have 64 dimensions per head. Additionally, Virchow 2G adopts additional normalization via DPN and QKN and has 8 registers. Virchow 2G is trained without solarization on a cosine schedule and is optimized with StableAdamW for a base learning rate equal to $1\times 10^{-4}$ and a batch size of $3072$ over a total of 2B balanced tiles. For stability, the second moment optimizer hyperparameter was reduced from $0.999$ to $0.95$. Additionally, both iBOT and DINO head teacher temperatures are reduced from $0.07$ to $0.04$. Further configuration details are found in appendix A3.
@@ 6.3 Embedding generation
For a 224 × 224 input tile image, a Virchow embedding is defined as the concatenation of the class token and the mean across all 256 of the patch tokens. This produces an embedding size of 2,560 (1,280 $\times$ 2) for Virchow, or 3,584 (1,792 $\times$ 2) for Virchow 2 / 2G. The class token is used for H-optimus-0 [44], Prov-GigaPath [57], UNI [12], and Phikon [24]. For CTransPath [53], the mean of all tokens is used as there is no class token.
@@ 6.5 Scaling results
Virchow 2 and 2G, along with Virchow and other baseline foundation models, were evaluated on in-distribution (ID) tile-level benchmarks (Tab. 2) and out-of-distribution (OOD) tile-level benchmarks (Tab. 3). The pathology-specific modifications to the training recipe of Virchow 2, along with increased training data scale and diversity, yielded a substantial boost in ID performance, raising the average weighted F1 score from 0.944 (Virchow) to 0.966 (Virchow 2). Gains in ID performance from further scaling the model size with Virchow 2G were more modest, further raising the score to 0.971. OOD average weighted F1 score increased from 0.877 with Virchow to 0.885 with Virchow 2 and 0.894 with Virchow 2G, demonstrating the benefits of model and data scale, as well as algorithmic improvements. Indeed, Fig. 1 demonstrates a log-linear trend in average OOD performance with model and data scale, across all models.
Although H-optimus-0 and Prov-GigaPath both use a ViT-g with 1.1B parameters, Virchow with ViT-H and 632M parameters matches or exceeds their performance on average, being surpassed only on PanMSK 20×20$\times$20 ×, MHIST, PCam, and WILDS by H-optimus-0 and on PanMSK 5×5$\times$5 ×, PCam, and WILDS by Prov-GigaPath. Virchow 2, also with 632M parameters but with a larger training set and improved training recipe, outperforms H-optimus-0 and Prov-Gigapath on all ID (PanMSK) tasks and MHIST, and bridges the gap on WILDS. Scaling up the model size to ViT-G (1.8B parameters) with Virchow 2G, results in top performance on all benchmark tasks. We note that while Virchow 2G achieves a weighted F1 score of 0.948 on TILS, as compared to the top score of 0.949 (Virchow, H-optimus-0), these scores are similar and performance on this task appears to have saturated. A major source of improvement lies in the addition of mixed magnifications. Virchow 2 and Virchow 2G both significantly improve on ID PanMSK tasks at 10×10$\times$10 × and 5×5$\times$5 × magnification. High resolution 40×40$\times$40 × OOD tasks such as MIDOG and DLBCL further illustrate the benefits of training on magnifications other than 20×20$\times$20 ×.
@@
Tab. 4 presents the Pearson correlation coefficient results for the HEST-benchmark. Here we observe less consistency in the performance of the models, possibly because the training protocol is fixed and the hyperparameters are not tuned. Virchow 2G has the best performance on 1/10 tasks, second best on 5/10 tasks, and the best average performance. H-optimus-0 has the best performance on 3/10 tasks, second best on 1/10 tasks, and the second best average performance.
@@
Table 2: Weighted F1 score results for the in-distribution tile-level benchmark tasks. Best results are bolded and second best are in italics. For the Virchow models, embeddings are defined as the concatentation of class and mean patch tokens. [CLS] indicates the use of class tokens only.
@@
Table 3: Weighted F1 score results for the out-of-distribution tile-level benchmark tasks. Best results are bolded and second best results are in italics. For the Virchow models, embeddings are defined as the concatentation of class and mean patch tokens. [CLS] indicates the use of class tokens only.
@@
Table 4: Pearson correlation results for the HEST-Benchmark tasks. Best results are bolded and second best results are in italics. For the Virchow models, embeddings are defined as the concatentation of class and mean patch tokens. [CLS] indicates the use of class tokens only.
@@ 7 Discussion
Further adjustments to the training recipe were required to allow scaling Virchow both along the data size and the model size axes. Without KDE, DINOv2 training collapsed before enough of the 3.1M slide dataset could be processed. Additional stabilizations were required to then scale up the model from ViT-H (632M parameters) to ViT-G (1.8B parameters). We note that largest models previously trained with DINOv2 had 1.1B parameters (ViT-g) [42, 44, 57]. The types of challenges encountered are classified into numerical issues and collapse events. Numerical issues are determined when a training run fails due to NaNs that could not be caught and mitigated and may be due to the use of FP16 rather than BF16. These are addressed by DPN, QKN, and gradient and update clipping in StableAdamW. Collapse events are categorized by sudden spikes in a loss component when features collapse to a lower dimensional subspace that is correlated with the degradation in downstream performance. Collapse primarily occurs in the iBOT head and is mitigated by reducing the learning rate, reducing the $\beta 2$ momentum parameter in StableAdamW from $0.999$ to $0.95$, and reducing the teacher temperature from $0.07$ to $0.04$ based on the guidance in DINO [11]. No collapse was observed in the DINO head. It is possible that diversity regularization provided sufficient support to avoid catastrophic events where asymmetric centering and sharpening was insufficient. These modifications reduce the convergence rate but allow training longer.
@@ A.1 Foundation models in CPath
Virchow 2G MSKCC 3.4M 1.8B ViT-G 1.8B DINOv2
Virchow 2 MSKCC 3.4M 1.7B ViT-H 632M DINOv2
@@ A.2 Hyperparameters
Augmentations Tile Context Size 392
@@
Augmentations Tile Context Size 392
@@
Optimizer Epsilon $1\times 10^{-8}$

### v2

@@ Abstract
Foundation models are rapidly being developed for computational pathology applications. However, it remains an open question which factors are most important for downstream performance with data scale and diversity, model size, and training algorithm all playing a role. In this work, we propose algorithmic modifications, tailored for pathology, and we present the result of scaling both data and model size, surpassing previous studies in both dimensions. We introduce two new models: Virchow2, a 632 million parameter vision transformer, and Virchow2G, a 1.9 billion parameter vision transformer, each trained with 3.1 million histopathology whole slide images, with diverse tissues, originating institutions, and stains. We achieve state of the art performance on 12 tile-level tasks, as compared to the top performing competing models. Our results suggest that data diversity and domain-specific methods can outperform models that only scale in the number of parameters, but, on average, performance benefits from the combination of domain-specific methods, data scale, and model scale.
footnotetext: These authors jointly supervised the work
@@ 1 Introduction
Figure 1: Virchow2 and Virchow2G adopt a pathology-adapted training method, scale up model and data size, and increase data diversity as compared to Virchow. Gains in the average weighted F1 score across 8 tile-level benchmarks using public data (see Tab. 3) are shown with respect to two scaling axes: model size (left), in terms of the number of parameters, and dataset size (right), in terms of the number of whole slide images (WSIs). Virchow2 improves over Virchow with pathology-specific modifications to the training method while scaling up the dataset from 1.5M to 3.1M WSIs, with increased diversity. Virchow2G then scales the model size from 632M to 1.9B parameters (ViT-H to ViT-G). Benchmark performance across foundation models in pathology appears to scale with model and dataset size.
In this work, we explore scaling along both the data and model size axes, while adjusting the training method with domain-inspired modifications (Fig. 1). We present Virchow2 and Virchow2G, mixed magnification foundation models for computational pathology. Virchow2, a 632M parameter vision transformer (ViT-H) expands on our work developing Virchow [^1] Virchow model weights are available for download at https://huggingface.co/paige-ai/Virchow [51], scaling the dataset size from 1.5M WSIs from a single institution with routine hemotoxylin and eosin (H&E) staining, to 3.1M WSIs from globally diverse institutions with diverse staining such as immunohistochemistry (IHC) [^2] Virchow2 model weights are available for download at https://huggingface.co/paige-ai/Virchow2. Virchow2G builds on Virchow2 by increasing the model size to 1.9B parameters (ViT-G).
@@ 2.2 CPath Foundation Models
3 Adapting self-supervised learning to the pathology domain
@@ 3.1 Domain-specific augmentations
Figure 2: Highlighted modifications of Virchow tailored for pathology. a. Virchow2 and Virchow2G are trained with mixed magnification samples; slides are tiled at 40x, 20x, 10x, or 5x. b. Illustration of the difference between standard crop-and-resize augmentation (top) and the proposed extended-context translation (ECT) augmentation (bottom). ECT is expected to be more suitable for pathology images as it avoids distorting the cell morphology. We tuned the ECT approach to retain the average overlap between crops (shaded, teal). c. A diversity regularization objective seeks to push apart tile embeddings on a hypersphere (top). For Virchow2 and Virchow2G, we replace the unstable KoLeo regularizer used by Virchow with a kernel density estimator (KDE) that does not produce exploding gradients for similar tiles (bottom), which is particularly important for pathology whole slide images as they contain many highly correlated tiles.
@@ 4 Stabilizing self-supervised vision transformers at scale
Optimizing ViTs requires adaptive methods like AdamW and its variations [5]. The stability of AdamW depends on the efficacy of the warm-up phase, on the learning rate, and on the selection of momentum parameters. Gradient spikes have been correlated with the ratio of first and second moments for an update step which provide insights on when a collapse event is likely to occur [39, 57]. For example, instability has been observed in the weights of the patch embedding layer [17, 57]. It is possible to reduce these issues by lowering the second moment hyperparameter but this may incur reduced downstream task performance. StableAdamW is explored as a means of stabilizing weight updates when the ratio between the first and second moments are too large (such as when nearly converged weights suddenly need a correction during a long training run) by mitigating large update steps with root-mean-square clipping [57].
@@ 5.3 Tile benchmarks
PanMSK is an in-distribution cancer detection task performed at 5$\times$, 10$\times$, and 20$\times$ magnifications. 1,196,171 $224\times 224$ pixel tile samples, sourced from 3,999 WSI across 2534 patients from MSKCC representing 17 tissue types, are labeled either cancer or benign [51]. To ensure a sufficient sample size for evaluation, no tissue (except for endometrial) has fewer than 44 WSIs with cancer from no fewer than 25 patients (max 139, median 43).
@@ 5.5 Ablation results
Table 1: Weighted F1 results for seven evaluation tasks and five model settings with a ViT-B/16. For each change, a checkmark indicates a deviation from the standard DINOv2 training protocol. In-distribution evaluation tasks include PanMSK XX, while out-of-distribution tasks include PCam, CRC, MHIST, and MIDOG. Boldface indicates the best performance and an underline indicates the second best performance. In a majority of tasks and on average, ECT, KDE, and no solarization regularization performs best.
6 Results of large-scale training
@@ 6.1 Training data
Figure 3: Overview of the training data. a. A breakdown of the scale of the training dataset with 3.1M whole slide images (WSIs), showing the distribution of slides with hemotoxylin and eosin (H&E) and immunohistochemical (IHC) stains. b. The data are globally diverse, with 15% of the slides and 57% of the patients sourced from many institutions across all continents. These slides were submitted to MSKCC for review. c. The distribution of most severe diagnoses of the WSIs, in decreasing order of severity: cancer, precursor, benign, unknown. d. The top identified tissue types shown as a fraction of the WSIs. There are over 150 additional tissue types in the long tail, summarized here as ‘Other Tissues’ (shown gray, along with ‘Unknown Tissue’, on the right). Online data balancing when training Virchow2G affected these tissue distributions as shown by the magenta dotted line.
@@ 6.2 Pretraining methodology
Following the modifications made to the ViT architecture to maximize compute efficiency [43], we introduce Virchow2G, a ViT-G/14 that adopts similar architectural changes. Virchow2G increases the embedding dimension from 1,664 to 1,792 to be a multiple of 256 and increases the number of attention heads from 16 to 28 in order to have 64 dimensions per head. Additionally, Virchow2G adopts additional normalization via DPN and QKN and has 8 registers. Virchow2G is trained without solarization on a cosine schedule and is optimized with StableAdamW for a base learning rate equal to $1\times 10^{-4}$ and a batch size of $3072$ over a total of 2 billion balanced tiles. For stability, the second moment optimizer hyperparameter was reduced from $0.999$ to $0.95$. Additionally, both iBOT and DINO head teacher temperatures are reduced from $0.07$ to $0.04$. Further configuration details are found in appendix A3.
@@ 6.3 Embedding generation
For a 224 × 224 input tile image, a Virchow embedding is defined as the concatenation of the class token and the mean across all 256 of the patch tokens. This produces an embedding size of 2,560 (1,280 $\times$ 2) for Virchow, or 3,584 (1,792 $\times$ 2) for Virchow2 and Virchow2G. The class token is used for H-optimus-0 [45], Prov-GigaPath [58], UNI [13], and Phikon [25]. For CTransPath [54], the mean of all tokens is used as there is no class token.
@@ 6.5 Scaling results
Virchow2 and Virchow2G, along with Virchow and other baseline foundation models, were evaluated on in-distribution (ID) tile-level benchmarks (Tab. 2) and out-of-distribution (OOD) tile-level benchmarks (Tab. 3). The pathology-specific modifications to the training recipe of Virchow2, along with increased training data scale and diversity, yielded a substantial boost in ID performance, raising the average weighted F1 score from 0.944 (Virchow) to 0.966 (Virchow2). Gains in ID performance from further scaling the model size with Virchow2G were more modest, further raising the score to 0.971. OOD average weighted F1 score increased from 0.877 with Virchow to 0.885 with Virchow2 and 0.894 with Virchow2G, demonstrating the benefits of model and data scale, as well as algorithmic improvements. Indeed, Fig. 1 demonstrates a log-linear trend in average OOD performance with model and data scale, across all models.
Although H-optimus-0 and Prov-GigaPath both use a ViT-g with 1.1B parameters, Virchow with ViT-H and 632M parameters matches or exceeds their performance on average, being surpassed only on PanMSK 20×20$\times$20 ×, MHIST, PCam, and WILDS by H-optimus-0 and on PanMSK 5×5$\times$5 ×, PCam, and WILDS by Prov-GigaPath. Virchow2, also with 632M parameters but with a larger training set and improved training recipe, outperforms H-optimus-0 and Prov-Gigapath on all ID (PanMSK) tasks and MHIST, and bridges the gap on WILDS. Scaling up the model size to ViT-G (1.9B parameters) with Virchow2G, results in top performance on all benchmark tasks. We note that while Virchow2G achieves a weighted F1 score of 0.948 on TILS, as compared to the top score of 0.949 (Virchow, H-optimus-0), these scores are similar and performance on this task appears to have saturated. A major source of improvement lies in the addition of mixed magnifications. Virchow2 and Virchow2G both significantly improve on ID PanMSK tasks at 10×10$\times$10 × and 5×5$\times$5 × magnification. High resolution 40×40$\times$40 × OOD tasks such as MIDOG and DLBCL further illustrate the benefits of training on magnifications other than 20×20$\times$20 ×.
@@
Tab. 4 presents the Pearson correlation coefficient results for the HEST-Benchmark. Here we observe less consistency in the performance of the models, possibly because the official training protocol uses fixed hyperparameters for all models’ embeddings, gene expression signal may not always be evident in H&E, and low sample counts for some tasks may add noise to the evaluation. However, the average performance across all ten gene expression prediction tasks scales with model size, as shown in Fig. A1c (Appendix A.3). Virchow2G (1.9B parameters) has the best average performance at 0.350 (best on 1/10 tasks, second best on 3/10 tasks), followed by H-optimus-0 (1.1B parameters) at 0.347 (best on 3/10 tasks, second best on 1/10 tasks). Results for all models are slightly higher in our evaluation than in the benchmark literature [31], possibly due to differences in supporting software library versions or non-determinism in the evaluation code.
@@
Table 2: Weighted F1 score results for the in-distribution tile-level benchmark tasks. Best results are bolded and second best are underlined. For the Virchow models, embeddings are defined as the concatentation of class and mean patch tokens. [CLS] indicates the use of class tokens only.
@@
Table 3: Weighted F1 score results for the out-of-distribution tile-level benchmark tasks. Best results are bolded and second best results are underlined. For the Virchow models, embeddings are defined as the concatentation of class and mean patch tokens. [CLS] indicates the use of class tokens only.
@@
Table 4: Pearson correlation results for the HEST-Benchmark tasks. Best results are bolded and second best results are underlined. For the Virchow models, embeddings are defined as the concatentation of class and mean patch tokens. [CLS] indicates the use of class tokens only. “Avg.” denotes the average correlation across all 10 tasks.
@@ 7 Discussion
Further adjustments to the training recipe were required to allow scaling Virchow both along the data size and the model size axes. Without KDE, DINOv2 training collapsed before enough of the 3.1M slide dataset could be processed. Additional stabilizations were required to then scale up the model from ViT-H (632M parameters) to ViT-G (1.9B parameters). We note that largest models previously trained with DINOv2 had 1.1B parameters (ViT-g) [43, 45, 58]. The types of challenges encountered are classified into numerical issues and collapse events. Numerical issues are determined when a training run fails due to NaNs that could not be caught and mitigated and may be due to the use of FP16 rather than BF16. These are addressed by DPN, QKN, and gradient and update clipping in StableAdamW. Collapse events are categorized by sudden spikes in a loss component when features collapse to a lower dimensional subspace that is correlated with the degradation in downstream performance. Collapse primarily occurs in the iBOT head and is mitigated by reducing the learning rate, reducing the $\beta 2$ momentum parameter in StableAdamW from $0.999$ to $0.95$, and reducing the teacher temperature from $0.07$ to $0.04$ based on the guidance in DINO [12]. No collapse was observed in the DINO head. It is possible that diversity regularization provided sufficient support to avoid catastrophic events where asymmetric centering and sharpening was insufficient. These modifications reduce the convergence rate but allow training longer.
@@ A.1 Foundation models in CPath
Virchow2G MSKCC 3.1M 1.9B ViT-G 1.9B DINOv2
Virchow2 MSKCC 3.1M 1.7B ViT-H 632M DINOv2
@@ A.2 Hyperparameters
Augmentations
Tile context size 392
@@
Augmentations
Tile context size 392
@@
Optimizer epsilon $1\times 10^{-6}$
@@
A.3 Scaling plots
Figure A1: The average performance across in-distribution tasks (a) and out-of-distribution tasks (b, c) scales with model size (left) and training data size (right). In-distribution data is sourced from the same institutions as Virchow training data, without any patient-level overlap.